# Time Series Forecasting with LLMs: Understanding and Enhancing Model Capabilities

### v3

@@ 5.1 External Knowledge Enhancing Time Series Forecasting
We introduce a novel method to improve the performance of large language models for time series forecasting. The core idea of this part is to use the knowledge obtained from the pre-training stage to help predict. We provide the large language model with some basic information about the current dataset such as the background of the data collection, and this process does not involve data leakage. We incorporate our tests on the data leakage in Appendix A.4. It is noted that we do not provide the LLMs with any statistical information such as the periods or trends. This approach ensures that the LLMs forecast the time series entirely based on the data and their prior knowledge. Let denote the initial prompt representing the original time sequence, and let denote the additional information. Consequently, the new prompt can be expressed as: .
@@ 5.1.1 Implementation Details
We input the dataset’s external knowledge through prompts before the sequence’s input. The external knowledge of each dataset is presented in subsection A.1. The results are shown in Table 8, where LLMTime Prediction refers to the approach described by [10] without any modifications.
@@ 5.1.2 Key Findings
5.2 Natural Language Paraphrasing
@@
We use natural language to describe the trend between consecutive values. For instance, given a time series where , we describe the trend from to as follows: "The value rises from to , and falls from to …". The string we get here is our natural language paraphrasing sequence. After generating responses based on the string, we extract the values from the text and construct the predicted time series.
@@ 6.1 Traditional Time Series Forecasting
Two commonly used methods for traditional time series analysis are the ARIMA method [2] and the exponential smoothing method [7]. The ARIMA model is a classic forecasting method that breaks down a time series into auto-regressive (AR), difference (I), and moving average (MA) components to make predictions. On the other hand, exponential smoothing is a straightforward yet effective technique that forecasts future values by taking a weighted average of past observations. The ARIMA model requires testing data stationarity and selecting the right order. However, the exponential smoothing method is not affected by outliers, it is only suitable for stationary time series, and its accuracy in predicting future values is lower than the ARIMA model.
@@ 6.2 LLMs for Time Series Forecasting
The first family of methods involve either pre-training a foundational large language model or fine-tuning existing LLMs by leveraging extensive time-series data [17, 8, 6, 3]. For instance, [17] aimed to build the foundational models for time series and investigate its scaling behavior. [4] proposed a two-stage fine-tuning strategy for handling multivariate time-series forecasting. Although these studies contribute significantly to understanding foundational models, they require considerable computing resources and expertise in fine-tuning procedures. Moreover, the details of the model may not be disclosed for commercial purposes [8], which impedes future research. Additionally, in scenarios with limited data available, there is insufficient information for training or fine-tuning.
In contrast, the second family of methods does not involve model parameter finetuning. These methods either create appropriate prompts or reprogramme inputs, to effectively handle time series data [10, 19, 14, 21]. [19] tokenizes the time series and manages to embed those tokens, and [14] reprogrammed the time series data with text prototypes before feeding them to the LLMs. These studies illuminate the characteristics of time series data and devise methods to align them with LLMs. However, they lack an analysis of the ability and bias in forecasting time series. The most related work to us is [10], though it lacks a quantitative analysis of the preference for the time series in LLMs, and it fails to explore the impact of input forms and prompt contents, such as converting the numerical time series into the natural language sequences and incorporating the background information into the prompt. Our work fills the gap, and we expect our work to be the benchmark for time-series analysis and provide insights for subsequent research.
@@ 7 Conclusions and Future Work
In this work, we investigate the key preferences of LLMs in the domain of time series forecasting under the zero shot setting, revealing a proclivity for data with distinct trends and seasonal patterns. Through a blend of real and synthetic datasets, coupled with counterfactual experiments, we have demonstrated LLMs’ improved forecasting performance with time series that exhibit clear periodicity. Besides, our results indicate that LLMs struggle with multi-period time series datasets, as they face difficulty in recognizing the distinct periods within them. Our findings also suggest that large language models are more sensitive to the segment of input sequences closer to the last known data than other locations. Lastly, experimental results indicate that our proposed strategies of incorporating external knowledge and transforming numerical sequences into natural language formats have yielded substantial improvements in accuracy.
Limitation
This study may be limited in the following ways. First, limitations in the scope of the dataset and large language models may not capture the full variability of the results with a wider array. In addition, some experimental sessions lack a comparison with hard-coded solutions, and there is a gap in understanding the performance of LLMs compared to traditional programming methods. Furthermore, the inability to categorize datasets by type and conduct specific types of experiments limits insight into the model’s performance in different data domains. These limitations suggest that the results could benefit from more extensive experiments and more nuanced analyses, underscoring the need to expand future research.
Appendix A Appendix
A.1 Dataset description and the External Knowledge incorporated in the Prompts
In this subsection, we briefly introduce the datasets we use, which also serve as the external knowledge incorporated into the prompts. Following [10], we downsample the input series to an hourly frequency, yielding a total of 267 observations and resulting in relatively small datasets. Additionally, we incorporate Memorization datasets published after September 2021, the cutoff date for GPT-3.5-turbo, to demonstrate the effectiveness of TimeLLM and our proposed methods. Finally, we implemented univariate time series forecasting to predict the ’OT’ feature on the ETTh1 and ETTm2 datasets, focusing on the last 96 steps of the test set.
A.1.1 Darts [11]
(1)AirPassengersDataset: This is a series of monthly passenger numbers for international flights, where each value is in thousands of passengers for that month.
(2)AusBeerDataset: This is a quarterly series of beer production, with each value representing the kiloliters of beer produced in that quarter.
(3)GasRateCO2Dataset: This time series dataset describes monthly carbon dioxide emissions.
(4)MonthlyMilkDataset: This time-series data set describing monthly milk production. Each is the average number of tons of milk each cow produces during the month.
(5)SunspotsDataset: This dataset records the number of sunspots each month, where each data is the number of sunspots in that month.
(6)WineDataset: This is a dataset of monthly wine production in Australia, where each figure is the number of wine bottles produced in that month.
(7)WoolyDataset: This is an Australian yarn production for each quarter, where each value is how many tons of yarn were produced in that quarter.
(8)HeartRateDataset: The series contains 1800 uniformly spaced instantaneous heart rate measurements from a single subject.
(9)ETTh1: This is a time series dataset containing high-frequency energy data of a certain region in China, which is mainly used for energy load forecasting and related time series analysis research.
(10)ETTm2: Similar to ETTh1, ETTm2 is also an energy time series dataset, but the data frequency or region covered may be different, which is also used for the analysis and prediction of energy consumption.
A.1.2 Memorization Datasets [10]
(11)TSMCStockDataset: This is historical trading data about Taiwan Semiconductor Manufacturing Corporation (TSMC) stock, containing information such as share price, volume, and date, and is commonly used in financial analysis and stock market forecasting research.
(12)TurkeyPowerDataset: This is a time series dataset on national electricity consumption in Turkey, which records the electricity usage in Turkey over a period of time and is often used to analyze and forecast electricity demand.
(13)IstanbulTrafficDataset: This dataset offers hourly Traffic Index data for Istanbul from October 2022 to May 2023.
A.1.3 Monash Datasets [9]
(14)US Births Dataset: This dataset contains the number of births in the US from 1969 to 1988.
(15)Saugeen River Flow Dataset: This dataset contains the daily mean flow of the Saugeen River at Walkerton in cubic meters per second from 1915 to 1979.
A.2 Implementation Details
In this subsection, we provide a comprehensive overview of the experiments conducted to investigate the preferences of LLMs for input time series data. We first describe both the real and synthesized datasets we use and then detail the methods we use to investigate the preferences of LLMs.
A.2.1 Real Datasets
We begin by comparing the performance of Large Language Models (LLMs) with traditional time series forecasting methods. The results are depicted in Figures Figure 4 and Figure 5, while Tables Table 9 and Table 10 present the computed metrics in tabular format.
To understand the preferences of LLMs, we conducted experiments on ten commonly used datasets: HeartRateDataset, GasRateCO2Dataset, AirPassengersDataset, AusBeerDataset, MonthlyMilkDataset, SunspotsDataset, WineDataset, WoolyDataset, IstanbulTrafficDataset and TurkeyPowerDataset. We apply the Seasonal-Trend decomposition using the LOESS (STL) technique to decompose the original time series into trend, seasonal, and residual components. In those datasets, we obtain the periods through the nature of the data. For instance, the number of passengers is collected monthly in AirPassengersDataset, and it’s natural to obtain that the period is 12. For the datasets without explicit periods, such as the IstanbulTrafficDataset, the period is determined through the periodogram, a widely used tool in signal processing assisting the identification of the time series period. The strengths and the model performance can be seen in Table 6.
Subsequently, we compute trend strength and seasonal strength to measure all of those components. We use MAPE to compute the Pearson correlation coefficients (PCC) across every two indexes and observe a relatively strong correlation between the strengths and model performance, signifying that LLMs perform better when the input time series possesses higher trend and seasonal strength (Shown in Table 1). Notably, GPT-4-turbo achieved a higher absolute PCC compared to GPT-3.5-turbo. It may be attributed to human feedback during GPT-4-turbo training, as individuals may be more aware of seasonal and trend data. This may provide some insights for further research into the characteristics of the LLMs with time series forecasting.
We also conduct a counterfactual analysis using a systematic permutation of the input time series. We begin by defining a sliding window that determines the length of the periods and add Gaussian noise to the data within this window. To reduce costs, we move the window by the period length. This method allows us to assess the importance of segments that LLM predominantly focuses on. Our observations suggest that introducing noise towards the end of the time series significantly affects LLM’s performance, leading to the inference that LLM tends to give more weight to the latter part of the time series in most instances.
A.2.2 Synthesized Datasets
To investigate the influence of the number of periods on model performance, we generated a dataset using the function , where denote the coefficients of the trend and seasonal components. We set , , and , uniformly sampled for 10 instances each, and and to represent the chosen frequencies. Similar to the previous experiments, ranges from 0 to 20 and follows the normal distribution . Our results reveal that LLMs exhibit worse performance when input sequences contain multiple periods, even when the seasonal strength is carefully controlled to be nearly unchanged, as is shown in Figure 8. This observation may be attributed to the LLMs’ challenge in recognizing and adapting to multiple periods, similar to human behavior.
A.3 Detailed Prompts for the Period Prediction
GPT-3.5-turbo and GPT-4-turbo.
You are a helpful assistant that specializes in time series analysis. The user will provide a sequence. The sequence is separated by commas. You need to infer the most probable underlying cycle of the sequence, even though there may also be a trend in this sequence. Do not assume that the underlying cycle has to stick to the popular cycles like 7 (days in a week), 12 (months in a year) and 30 (days in a month), just infer the cycle fully based on the inherent cycle of the given sequence. The underlying cycle of the sequence is strictly below 8. Please infer the cycle without producing any additional text. Again, the sequence is separated by commas. Sequence:
Gemini.
You are a helpful assistant that specializes in time series analysis. The user will provide a sequence. The sequence is represented by decimal strings separated by commas. You need to infer the most probable underlying cycle of the sequence, even though there may also be a trend in this sequence. Do not assume that the underlying cycle has to stick to the popular cycles like 7 (days in a week), 12 (months in a year) and 30 (days in a month), just infer the cycle fully based on the inherent cycle of the given sequence. The underlying cycle of the sequence is strictly below 8. Please infer the cycle without producing any additional text. Sequence:
A.4 Tests on Data Leakage
Since the detailed datasets used for training GPT-4 and Gemini are not directly accessible, we conducted several experiments to indirectly investigate the data leakage issue. First, we performed an "Acknowledge Test" where we asked the LLMs if they recognize the dataset based on its name. Next, we conducted a "Series Test" by asking the LLMs to predict the first 20 steps of the datasets. Finally, we carried out a "Dataset Detection" test, where we fed the first 20 steps into the LLMs and asked them to identify the dataset. All these experiments help us determine whether the LLMs only recognize the name or have detailed knowledge of the datasets.
As shown in 4, we observe that both GPT and Gemini are generally aware of the content of most datasets. However, detailed sequence data is primarily known only for the AirPassengers dataset when using GPT, and this level of detail is not known for other datasets. During the experiments on the Dataset Detection, it is found that both GPT and Gemini could identify which dataset is being used based on the first 20 data points of the time series. All these results indicate that while they can recognize and identify datasets from limited information, they typically do not possess detailed knowledge of the sequence data for a broader range of datasets.
@@ A.4 Tests on Data Leakage
A.5 Computational Cost
We list the average token length cost associated with external knowledge enhancing and natural language paraphrasing for reference. Avg Token Length(ori) is the prompt Length of the unexecuted method, and Avg Token Length(EKE, NLP) is the prompt length after executing the corresponding policy. It is noted that Natural Language Paraphrasing is judged one by one through hard coding. Besides, there is a length check after transformation, so it is guaranteed that a certain length can be obtained each time. The results are shown in 5.

### v4

@@ 5 How to Leverage These Insights to Improve the Model’s Performance?
5.1 Dataset description and the External Knowledge incorporated in the Prompts
We briefly introduce the datasets we use, which also serve as the external knowledge incorporated into the prompts. Following [5], we downsample the input series to an hourly frequency, yielding a total of 267 observations and resulting in relatively small datasets. Additionally, we incorporate Memorization datasets published after September 2021, the cutoff date for GPT-3.5-turbo, to demonstrate the effectiveness of TimeLLM and our proposed methods. Finally, we implemented univariate time series forecasting to predict the ’OT’ feature on the ETTh1 and ETTm2 datasets, focusing on the last 96 steps of the test set.
@@ 5.2 External Knowledge Enhancing Time Series Forecasting
We introduce a novel method to improve the performance of large language models for time series forecasting. The core idea of this part is to use the knowledge obtained from the pre-training stage to help predict. We provide the large language model with some basic information about the current dataset, such as the background of the data collection, and this process does not involve data leakage. We incorporate our tests on the data leakage in Section 5.5. We do not provide the LLMs with any statistical information, such as the periods or trends. This approach ensures that the LLMs forecast the time series entirely based on the data and their prior knowledge. Let denote the initial prompt representing the original time sequence, and let denote the additional information. Consequently, the new prompt can be expressed as: .
@@ 5.2.1 Implementation Details
We input the dataset’s external knowledge through prompts before the sequence’s input. The external knowledge of each dataset is presented in subsection 5.1. The results are shown in Table 8, where LLMTime Prediction refers to the approach described by [5] without any modifications.
@@ 5.2.2 Key Findings
5.3 Natural Language Paraphrasing Time Series Forecasting
@@
We use natural language to describe the trend between consecutive values. For instance, given a time series where , we describe the trend from to as follows: ”The value rises from to , and falls from to …”. The string we get here is our natural language paraphrasing sequence. After generating responses based on the string, we extract the values from the text and construct the predicted time series.
@@ 5.3.1 Implementation Details
We use GPT-3.5-Turbo, GPT-4-turbo, Llama-2 and Gemini-Pro-1.0 to forecast the time series, where part of the results are presented in Table 4 due to the page limit.
@@ 5.3.2 Key Findings
5.4 Computational Cost
For reference, we list the average token length cost associated with external knowledge enhancement and natural language paraphrasing. Avg Token Length(ori) is the prompt Length of the unexecuted method, and Avg Token Length(EKE, NLP) is the prompt length after executing the corresponding policy. It is noted that Natural Language Paraphrasing is judged one by one through hard coding. Besides, there is a length check after transformation, so it is guaranteed that a certain length can be obtained each time. The results are shown in Table 7. Several key observations can be made: the original TimeLLM method maintains a uniform token length of 200 across all datasets, providing a stable baseline. EKE results in a slight increase in token length, ranging from 7% to 12%, suggesting a good balance between incorporating additional context and maintaining computational efficiency. In contrast, NLP leads to a more substantial increase in token numbers.
5.5 Tests on Data Leakage
We indirectly explore the data leakage problem by asking LLMs if they can identify the dataset name, the first 20 steps of the predicted dataset, and identify the dataset based on the first 20 steps of the time series data points. The results show that although GPT and Gemini can identify and determine data sets with limited information, they generally do not have detailed sequence data knowledge for a wider range of data sets Table 6.
@@ 5.5 Tests on Data Leakage
| MSE | MAE | MAPE | MSE | MAE | MAPE | ||
| AirPassengers | 3713.99 | 50.37 | 10.88 | 6244.07 | 61.39 | 14.43 | |
| AusBeer | 669.01 | 21.82 | 5.12 | 841.68 | 23.59 | 5.62 | |
| GasRateCO2 | 16.47 | 3.36 | 5.97 | 10.88 | 2.66 | 4.73 | |
| MonthlyMilk | 4781.26 | 55.45 | 6.25 | 7507.13 | 66.28 | 112.77 | |
| GPT-3.5-turbo-1106 | Sunspots | 7072.42 | 62.61 | 194.29 | 6556.55 | 58.95 | 217.94 |
| HeartRate | 59.83 | 6.44 | 6.75 | 76.83 | 7.15 | 7.42 | |
| Istanbul-Traffic | 888.31 | 28.16 | 60.11 | 1321.44 | 48.7 | 7.47 | |
| ETTh1 | 2.65 | 1.01 | 132.13 | 5.64 | 2.71 | 1.625 | |
| ETTm2 | 2.00 | 0.89 | 201.84 | 3.46 | 2.17 | 1.178 | |
| AirPassengers | 1262.24 | 30.54 | 6.80 | 1286.25 | 28.04 | 6.07 | |
| AusBeer | 345.59 | 15.70 | 3.69 | 513.49 | 18.57 | 4.28 | |
| GasRateCO2 | 6.99 | 2.29 | 4.21 | 7.27 | 2.32 | 4.18 | |
| MonthlyMilk | 2209.33 | 44.02 | 5.12 | 4442.18 | 50.75 | 172.82 | |
| GPT-4-turbo-preview | Sunspots | 4571.92 | 50.24 | 334.30 | 3374.70 | 41.87 | 321.11 |
| HeartRate | 78.99 | 6.96 | 7.90 | 988.14 | 26.57 | 29.22 | |
| Istanbul-Traffic | 954.88 | 26.92 | 47.29 | 1291.17 | 32.16 | 6.46 | |
| ETTh1 | 2.70 | 1.06 | 129.99 | 4.73 | 1.53 | 3.282 | |
| ETTm2 | 1.18 | 0.79 | 291.67 | 2.30 | 1.034 | 1.607 | |
| AirPassengers | 3713.99 | 50.37 | 10.88 | 1286.25 | 28.04 | 6.07 | |
| AusBeer | 893.56 | 21.49 | 4.87 | 644.82 | 17.88 | 4.08 | |
| GasRateCO2 | 11.38 | 3.04 | 5.49 | 12.78 | 2.97 | 5.47 | |
| MonthlyMilk | 4722.32 | 60.36 | 7.05 | 3410.20 | 41.40 | 240.25 | |
| Llama-2 | Sunspots | 4000.19 | 46.45 | 138.69 | 4467.67 | 48.95 | 91.79 |
| HeartRate | 112.17 | 7.86 | 8.93 | 75.58 | 7.11 | 7.94 | |
| Istanbul-Traffic | 979.15 | 26.70 | 45.57 | 1531.37 | 34.74 | 7.42 | |
| ETTh1 | 4.15 | 1.65 | 408.11 | 4.84 | 1.79 | 3.178 | |
| ETTm2 | 3.08 | 1.47 | 810.56 | 3.31 | 2.07 | 2.153 | |
| AirPassengers | 5237.85 | 51.92 | 11.08 | 6392.21 | 63.57 | 14.03 | |
| AusBeer | 325.45 | 10.84 | 1.86 | 397.78 | 14.36 | 3.27 | |
| GasRateCO2 | 15.54 | 3.23 | 4.43 | 18.99 | 3.57 | 6.46 | |
| MonthlyMilk | 491.26 | 15.18 | 1.13 | 628.98 | 17.01 | 1.99 | |
| Gemini-1.0-pro | Sunspots | 491.64 | 11.15 | 1.27 | 626.03 | 14.94 | 1.73 |
| HeartRate | 47.45 | 4.83 | 4.67 | 57.96 | 6.01 | 6.66 | |
| Istanbul-Traffic | 1253.74 | 28.25 | 5.42 | 1531.37 | 34.74 | 7.42 | |
| ETTh1 | 2.92 | 1.45 | 2.88 | 4.84 | 1.79 | 3.178 | |
| ETTm2 | 2.00 | 1.74 | 1.22 | 3.31 | 2.07 | 2.153 | |
@@ 6.1 Traditional Time Series Forecasting
Two commonly used traditional time series analysis methods are the ARIMA method [16] and the exponential smoothing method [17]. The ARIMA model is a classic forecasting method that breaks down a time series into auto-regressive (AR), difference (I), and moving average (MA) components to make predictions. On the other hand, exponential smoothing is a straightforward yet effective technique that forecasts future values by taking a weighted average of past observations. The ARIMA model requires testing data stationarity and selecting the right order. However, the exponential smoothing method is not affected by outliers; it is only suitable for stationary time series, and its accuracy in predicting future values is lower than that of the ARIMA model.
@@ 6.2 LLMs for Time Series Forecasting
The first family of methods involves either pre-training a foundational large language model or fine-tuning existing LLMs by leveraging extensive time-series data [6, 18, 19, 20]. For instance, [6] aimed to build the foundational models for time series and investigate its scaling behavior. [21] proposed a two-stage fine-tuning strategy for handling multivariate time-series forecasting. Although these studies contribute significantly to understanding foundational models, they require considerable computing resources and expertise in fine-tuning procedures. Moreover, the details of the model may not be disclosed for commercial purposes [18, 22], which impedes future research. Additionally, in scenarios with limited data available, there is insufficient information for training or fine-tuning.
In contrast, the second family of methods does not involve model parameter finetuning. These methods either create appropriate prompts or reprogramme inputs to effectively handle time series data [5, 7, 23, 24]. [7] tokenizes the time series and manages to embed those tokens, and [23] reprogrammed the time series data with text prototypes before feeding them to the LLMs. These studies illuminate the characteristics of time series data and devise methods to align them with LLMs. However, they lack an analysis of the ability and bias in forecasting time series. The most related work to us is [5], though it lacks a quantitative analysis of the preference for the time series in LLMs, and it fails to explore the impact of input forms and prompt contents, such as converting the numerical time series into the natural language sequences and incorporating the background information into the prompt. Our work fills the gap, and we expect our work to be the benchmark for time-series analysis and provide insights for subsequent research.
@@ 7 Conclusions and Future Work
In this work, we investigate the key preferences of LLMs in the domain of time series forecasting under the zero-shot setting, revealing a proclivity for data with distinct trends and seasonal patterns. Through a blend of real and synthetic datasets, coupled with counterfactual experiments, we have demonstrated LLMs’ improved forecasting performance with time series that exhibit clear periodicity. Besides, our results indicate that LLMs struggle with multi-period time series datasets, as they face difficulty in recognizing the distinct periods within them. Our findings also suggest that large language models are more sensitive to the segment of input sequences closer to the last known data than other locations. Lastly, experimental results indicate that our proposed strategies of incorporating external knowledge and transforming numerical sequences into natural language formats have yielded substantial improvements in accuracy.